### v1

# End-to-End Speech-to-Text Translation: A Survey

Abstract
Speech-to-text translation pertains to the task of converting speech signals in a language to text in another language. It finds its application in various domains, such as hands-free communication, dictation, video lecture transcription, and translation, to name a few. Automatic Speech Recognition (ASR), as well as Machine Translation(MT) models, play crucial roles in traditional ST translation, enabling the conversion of spoken language in its original form to written text and facilitating seamless cross-lingual communication. ASR recognizes spoken words, while MT translates the transcribed text into the target language. Such disintegrated models suffer from cascaded error propagation and high resource and training costs. As a result, researchers have been exploring end-to-end (E2E) models for ST translation. However, to our knowledge, there is no comprehensive review of existing works on E2E ST. The present survey, therefore, discusses the work in this direction. Our attempt has been to provide a comprehensive review of models employed, metrics, and datasets used for ST tasks, providing challenges and future research direction with new insights. We believe this review will be helpful to researchers working on various applications of ST models.
keywords:
Speech-to-Text Translation, Automatic Speech Recognition, Machine Translation, Modality Bridging
[…]
[label1]organization=Indian Institute of Technology Indore, country=India
1 Introduction
The Speech-to-Text (ST) translation task aims to convert a speech in one language into text in another language. It finds its applications in various areas such as automatic subtitling, dictations, video lecture translations, tourism, telephone conversations, to name a few. There are many facets under which the ST problem can be cast. For example, are we performing ST translation online (aka simultaneous translation) or offline? The former is required in live video streaming, while the latter is helpful for movies where some latency may be allowed. The ST problem is further exacerbated by noisy inputs, low-resource/code-mix languages, and the presence of multiple speakers.
The traditional ST translation methods involve a cascade approach– First, applying automatic speech recognition (ASR) on the given speech and then performing machine translation (MT) on the transcription produced by ASR. Such a cascade model is prone to several issues, such as (a) error in the ASR model can propagate to the MT model, (b) higher cost and training time, and (d) resources required for training. To mitigate such issues, various researchers propose using end-to-end (E2E) models for ST task (Bérard et al., 2016, 2018; Anastasopoulos et al., 2016; Bentivogli et al., 2021; Gangi et al., 2019). E2E models usually employ a single neural network that can be trained end-to-end. Because of simpler training, lower memory footprint, and cost, E2E model development has gained significant momentum in the research community. However, a systematic and comprehensive review of ST works is missing. Concurrently to our work, we find that a review paper (Xu et al., 2023) on ST was published recently. The aforementioned review categorizes existing works mainly based on modeling, data, and application issues. They do not cover the data sets available for the ST tasks nor provide any insights into the cascade vs. E2E model performances. Also, the future open problems provided by them are limited. On the other hand, our work comprehensively reviews the existing models for ST tasks and datasets from a completely different perspective and critically analyzes the existing works; thereafter, we identify several challenges and future research directions. Thus, our work may be deemed complimentary to (Xu et al., 2023).
The following review is structured as follows: In Section 2, we establish the foundation of the ST task through a formal definition, and we subsequently delve into the various metrics and loss functions adopted by different researchers. Moving on to Section 3, we present comprehensive descriptive statistics pertaining to benchmark datasets that have been employed by researchers within this field. Section 4 is dedicated to exploring the different toolkits accessible for addressing E2E ST tasks. In Section 5, we delve into the diverse strategies and approaches employed to tackle the ST problem. Initially, we categorize these approaches based on the frameworks utilized and the characteristics of the data involved. Under the framework categorization in Section 5.1, we provide explanations for both Sequence-to-sequence (seq2seq) frameworks (5.1.1) and Modality-bridging frameworks (5.1.2). Additionally, the seq2seq frameworks are further subcategorized, taking into account scenarios involving ST with MT, ST with ASR, and ST with both MT and ASR. Furthermore, the nature of the data is dissected into several distinctive settings in Section 5.2, including Low-resource (5.2.1), Code-mixed (5.2.2), Streaming (5.2.3), Unsupervised (5.2.4), and Multilingual (5.2.5). Finally, in Section 6, we explore the prospects for future research and address open problems within the field.
2 Background
This section describes the ST task formally and presents the metrics used to evaluate the performance of the ST model.
2.1 Task Definition
ST task can be defined as translating the given input speech $U$ in one language to translated text $V$ in another language with the transcription text $X$. The definition, in formal terms, is as follows: Given a dataset $D=\{({\bf u}^{i},{\bf x}^{i},{\bf v}^{i})|i=1,2,\ldots,n\}$ of pairs of input speech features ${\bf u}=(u_{1},u_{2},\ldots,u_{T_{u}})$ in a language and output text tokens ${\bf v}=(v_{1},v_{2},\ldots,v_{T_{v}})$ in a different language, the objective of the ST task is to minimize the conditional probability given below:
$p({\bf v}|{\bf u};\theta)=\prod_{t=1}^{T_{v}}p(v_{t}|v_{<t},{\bf u};\theta)$ (1)
In the above equation, $T_{u}$, $T_{v}$, and $\theta$ are the lengths of input features, the number of output tokens, and the model parameter, respectively. There are $n$ parallel sentences in our corpus, and the model is optimized for negative log-likelihood over these sentences as
$\ell(\theta|D)=-\sum_{i=1}^{n}\log P({\bf v}^{i}|{\bf u}^{i};\theta)$ (2)
The above optimization is usually solved using an encoder-decoder with an attention approach. Essentially, an encoder maps speech input to a hidden state representation $h$ followed by a decoder which takes the previously generated text tokens $v_{<t}$, encoder hidden state $h$ and attention vector ${\bf\alpha}$ (Vaswani et al., 2017). Offline ST translation can look at the whole speech before producing output text tokens, whereas online ST can start translation as soon as it sees a few seconds of the speech signal.
2.2 Evaluation Metrics
This section discusses various metrics used to evaluate the E2E ST models. The metrics of E2E ST models are categorized into 2 major performance measures which are quality and latency. The quality of the E2E ST models is the measure of how close the ST translation is to the target sentence or human-translated sentence. The latency of the E2E ST models is the time taken by the model to translate source speech into the target text.
Quality-based metrics: The quality-based metrics measure how close the translation is to the target sentence, which usually lies in [0, 1]. Most of the existing literature evaluates these scores on detokenized output. Standard translation metrics for evaluating ST task performance are presented below:
  • 1.
    BLEU Bi-lingual Evaluation Understudy (BLEU) Papineni et al. (2002) is the score of weighted average length of the translated sentence that matched the target sentence. The expression for evaluating BLEU score is given by:
    $BLEU=BP.exp(\sum_{n=1}^{N}w_{n}\log P_{n})$ (3)
    where the $BP$, $w_{n}$, and $P_{n}$ are brevity penalty, positive weights that sums to 1, and n-gram precisions of length upto $N$, respectively.
  • 2.
    TER Translation Edit Rate (TER) Snover et al. (2006) is the measure of how much a translated sentence has to be modified in order to match the target sentence. The expression for TER is given by:
    $TER=M/R$ (4)
    where $M$ and $R$ are the modifications required in the translated sentence and the average number of reference words, respectively.
  • 3.
    METEOR Metric for Evaluation of Translation with Explicit word Ordering (METEOR) Banerjee and Lavie (2005) is the automatic metric that matches the unigrams in the translated text to the referenced text. The expression for METEOR Score is given by:
    $M\_Score=F_{M}*(1-Penalty_{M}),$ (5)
    where $F_{M}=\frac{10PR}{R+9P},$ and $Penalty_{M}=0.5*\left(\frac{C}{U}\right)^{3}.$
    The $F_{M}$ and $Penalty_{M}$ are the F-Score and the penalty of the METEOR score. $P$, $R$, $C$, and $U$ denote the precision, recall, number of chunks, and number of unigrams matched in the translated text, respectively.
  • 4.
    BERTScore The BERTScore Zhang et al. (2019) is an automatic evaluation metric that scores the similarity between the translated text to the referenced text. It takes into account the Recall ($R$), Precision ($P$), and Fscore ($F$), which are expressed as:
    $R=\frac{1}{|v|}\sum_{v_{i}\in v}\max_{\hat{v_{j}}\in\hat{v}}v_{i}^{T}\hat{v_{j}},P=\frac{1}{|\hat{v}|}\sum_{\hat{v_{j}}\in\hat{v}}\max_{v_{i}\in v}v_{i}^{T}\hat{v_{j}},F=2.\frac{R.P}{R+P}$ (6)
    where $v_{i}$ and $\hat{v_{j}}$ are tokens of the sentences in the translated text and the ground truth text.
There are few other evaluation metrics such as TRANSTAC Schlenoff et al. (2009), NIST-Score Doddington (2002) and CHRF Popović (2015) which are less frequently used.
Latency-based metrics: For streaming ST task, researchers report a metric for measuring latency, which is defined as the delay incurred in starting to produce the translation. Let ${\bf u},{\bf v}$ and $\hat{{\bf v}}$ denote the input speech sequence, ground truth text sequence, and system-generated hypothesis sequence, respectively. In the streaming ST task, the model can only produce output with partial input. Suppose ${\bf u}_{1:t}=(u_{1},\ldots,u_{t}),t<T_{u}$ has been read when generating $v_{s}$, the delay in $v_{s}$ is defined as (Ma et al., 2020)
$d_{s}=\sum_{k=1}^{t}T_{k}$ (7)
where $T_{k}$ is the duration of the speech frame $u_{k}$. The latency metrics are evaluated using a method that analyzes a sequence of time delays $[d_{1},\ldots,d_{T_{v}}]$.
  • 1.
    Average Proportion (AP) calculates the mean fraction of the source input that is read during the target prediction generating process.
    $AP=\frac{1}{T_{y}\sum_{k=1}^{T_{u}}T_{k}}\sum_{s=1}^{T_{v}}d_{s}$ (8)
  • 2.
    Average Lagging (AL) measures the distance between the speaker and the user based on the number of words used in the conversation(Ma et al., 2018).
    $AL=\frac{1}{\tau(T_{u})}\sum_{s=1}^{\tau(T_{u})}d_{s}-\hat{d_{s}}$ (9)
    Where $\tau(T_{u})=\min\{s\mid d_{s}=\sum_{k=1}^{T_{u}}T_{k}\}$ and $\hat{d_{s}}$ are the delays of an ideal policy defined as (Ma et al., 2020)
    $\hat{d_{s}}=(s-1)\sum_{k=1}^{T_{u}}\frac{T_{k}}{\hat{T_{v}}}$ (10)
  • 3.
    Differentiable Average Lagging (DAL) One issue with AL is that it is not differentiable because of the $\min$ function. To solve this, (Cherry and Foster, 2019) introduces a minimum delay of $1/\gamma$ after each operation and defines DAL as
    $DAL=\frac{1}{T_{v}}\sum_{s=1}^{T_{v}}d_{s}^{{}^{\prime}}-\frac{s-1}{\gamma}$ (11)
    where
    $d_{s}^{{}^{\prime}}=\begin{cases}d_{s},&\ s=0\\ \max(d_{s},d^{{}^{\prime}}_{s-1}+\gamma),&s>0\end{cases}$ (12)
    and $\gamma=T_{v}/\sum_{k=1}^{T_{u}}T_{k}$
2.3 Loss Functions
This section discusses the various loss functions used for the E2E ST models. Let $D=(u,x,v)$ be a tuple where $u,x$, and $v$ are the speech, the transcription text, and the translation text, respectively. The following are the various loss functions that are used to optimize the performance of the E2E ST models:
  • 1.
    ST Loss (Ouyang et al., 2022) is defined as the negative log-likelihood of the translation text given the source speech as follows
    $L_{ST}=-\sum_{(u,v)\in D}\log p(v|u)$ (13)
  • 2.
    MT Loss (Ouyang et al., 2022) is defined as the negative log-likelihood of the translation text given the source transcript as follows
    $L_{MT}=-\sum_{(x,v)\in D}\log p(v|x)$ (14)
  • 3.
    ASR Loss (Ouyang et al., 2022) is defined as the negative log-likelihood of the transcription text given the source speech as follows
    $L_{MT}=-\sum_{(u,x)\in D}\log p(x|u)$ (15)
  • 4.
    CTC Loss(Ren et al., 2020) computes the most likely alignment of output text sequence given input speech sequence by summing over the all possible output sequence paths.
    $L_{CTC}=-\sum_{(u,x)\in D}\sum_{z\in\phi(x)}\log p(z|u)$ (16)
  • 5.
    Cross-Modal Adaptation Loss (Liu et al., 2020b) is defined as the sum of all the Mean Squared Errors of the speech and the transcription texts.
    $L_{AD}=\Biggl{\{}\begin{matrix}\sum_{(u,x)\in D}MSE(\bar{h_{u}},\bar{h_{x}});&&$seq-level$\\ \sum_{(u,x)\in D}MSE(h_{u},h_{x});&&$word-level$\end{matrix}$ (17)
    where $h_{u}$ and $h_{x}$ are the speech and word embeddings, and $\bar{h_{u}}$ and $\bar{h_{x}}$ are the average speech and word embeddings, respectively. MSE represents the difference between the two embeddings.
  • 6.
    Cross-Entropy Loss (Ye et al., 2021) is the negative likelihood of the data combined over all the subtasks such as ASR, MT, ST and also from external-MT.
    $L_{\theta}=-\sum_{x,v\in D^{\prime}\cup D_{MT-ext}}\log p(x|v;\theta),$ (18)
    where $D^{\prime}=D_{ASR}\cup D_{MT}\cup D_{ST}$ is the superset of all the parallel subsets data.
  • 7.
    Distillation Loss (Liu et al., 2019) The student model not only matches the ground truth, but also the teacher models’s output probabilities, which reduces the variance of the gradients.
    $L_{KD}=-\sum_{(x,v)\in D}\sum_{t=1}^{N}\sum_{k=1}^{|V|}S(v_{t}=k|v_{<t},x)\log T(v_{t}=k|v_{<t},x)$ (19)
    where $S$ and $T$ denote the output distribution of student and teacher models, respectively.
  • 8.
    Contrastive Loss (Ye et al., 2022a) It is computed between the the speech and the transcription text bringing them closer, and pushing the unrelated pairs farther.
    $L_{CON}=-\sum_{(u,x)\in D}\log\frac{\exp({\cos(\bar{h_{u}},\bar{h_{x}})}/\kappa)}{\sum_{\forall x_{j}\notin\bar{h_{x}}}\exp({\cos(\bar{h_{u}},\bar{h_{x}}(x_{j}))}/\kappa)},$ (20)
    where $cos$ and $\kappa$ denote the cosine similarity and temperature hyperparameter, respectively.
Datasets Source Language (Speech) Target Language (Text) Speech (hours) Speakers Validation Gender Age Group
MuST-C
En
14 lang
0.4K
1.6K
Librispeech
En
Fr
0.2K
1.4K
CoVost
En
11 lang
0.7K
11K
CoVost2
21 lang
En
2.8K
11K
En
15 lang
0.7K
78K
EuroparlST
4 lang
4 lang
0.25K
VoxPopuli
En
15 lang
1.79K
4.3K
Kosp2e
Ko
En
0.2K
0.2K
GigaST
En
De, Zh
10K
Prabhupadavani
en-bn-sn code-mix
25 lang
0.09K
0.13K
How2
En
Pt
2K
Table 1: Dataset statistics(✓ means that feature is available for the dataset and ✗ means that the feature is unavailable for the dataset)
3 Datasets for ST Tasks
There have been several datasets created for the ST task. Some of them are listed below, and we describe them here briefly. Table 1 provides information on various dataset statistics, such as hours of speech, the number of speakers, whether the dataset was manually or machine validated, the gender, and the age range to which the speaker belongs. Additionally, the tools required for creating these datasets are (a) Gentle (Ochshorn and Hawkins, 2017) for audio-transcription alignment, and (b) BertAlign [^1] https://github.com/bfsujason/bertalign for transcription-translation alignment.
  • 1.
    How2 (Sanabria et al., 2018) is an ST corpus of English instructional videos having Portuguese translations.
  • 2.
    Augmented Librispeech (Kocabiyikoglu et al., 2018) is obtained from the LibriSpeech corpus (Panayotov et al., 2015). It is a speech recognition repository generated using audiobooks of Gutenberg Project  [^2] https://www.gutenberg.org/. This dataset is designed to translate English speech into written French text.
  • 3.
    MuST-C (Di Gangi et al., 2019) It is a large multilingual ST translation corpus available . It contains translations from English into fourteen additional languages and is compiled from TED Talks. mTEDx (Salesky et al., 2021) is one such multilingual dataset from TED talks.
  • 4.
    CoVoST and CoVoST 2 (Wang et al., 2020a, c), the datasets are based on Common Voice project  [^3] https://commonvoice.mozilla.org/en. CoVoST is a many-to-one dataset covering 11 languages, while CoVoST 2 offers one-to-many and many-to-one translations for 15 languages.
  • 5.
    Europarl-ST (Iranzo-Sánchez et al., 2020) is a collection that contains speech and text data from European Parliament proceedings between 2008 and 2012 in four languages. It includes multiple sources and targets for both speech and text.
  • 6.
    VoxPopuli (Wang et al., 2021a) dataset is an expansion of Europarl-ST. It includes data from European parliament sessions spanning from 2009 to 2020.
  • 7.
    Kosp2e (Cho et al., 2021a) is a Korean (ko) to English(en) ST translation corpus, which contains Korean speech with parallel English texts. The corpus contains data from four different domains: Zeroth from news/newspaper, KSS (Park, 2018) from textbooks, emphStyleKQC (Cho et al., 2021b) from AI applications, and Covid-ED (Lee et al., 2021) from covid diaries of people which have emotions.
  • 8.
    GigaST (Ye et al., 2022b) corpus is a collection of speech translations from English to German and Chinese. It is created using the English ASR GigaSpeech(Chen et al., 2021a), which features 10,000 hours of transcribed speech from various sources such as audioPortugesebooks, podcasts, and YouTube.
  • 9.
    Prabhupadavani (Sandhan et al., 2022) is an ST dataset where speech is multilingual and Code-Mix with three different languages, English is the primary language, and words and phrases from Sanskrit and Bengali are interjected. The text part has sentences in 25 languages.
Besides these popular ST datasets, there are some other smaller size datasets such as Fisher(Cieri et al., 2004), Call-Home [^4] https://ca.talkbank.org/access/CallHome/eng.html, Gordard Corpus(Godard et al., 2017), Glosse Audio Corpus [^5] https://ainu.ninjal.ac.jp/folklore/en/, BTEC  [^6] http://universal.elra.info/product_info.php?cPath=37_39&products_id=80, WSJ [^7] https://catalog.ldc.upenn.edu/LDC93s6a, IWSLT [^8] https://iwslt.org/, CHiME-4 Corpus(Christensen et al., 2010), Miami Corpus(Deuchar, 2008), and MSLT Corpus (Federmann and Lewis, 2016).
4 Toolkits for ST
In order to facilitate building and training ST models, various researchers have proposed a few toolkits. The toolkits for ST create an environment where the dataset for ST tasks can be pre-processed, and models can be trained, fine-tuned, and evaluated. We provide a short description of these toolkits to make the survey a place for a one-stop-shop for ST modeling.
  • 1.
    EspNet-ST [^9] https://github.com/espnet/espnet toolkit(Inaguma et al., 2020) is developed as there was no toolkit available for performing the sub-tasks of ST. EspNet-ST provides ASR, LM, E2E-ST, Cascade-ST, MT, and TTS along with examples. It also provided pre-trained transformer-based models on various datasets like MUST-C, Libri-trans, Fisher, CALL-HOME, and How2.
  • 2.
    FairSeq S2T [^10] https://github.com/facebookresearch/fairseq/tree/main/examples/speech_to_text(Wang et al., 2020b) toolkit is an extension to FairSeq(Ott et al., 2019) in which all the functions of EspNet-ST are available. Additionally, it provides the Non-Autoregressive MT, Online ST, and Speech Pretraining. The toolkit also provides state-of-the-art ST models based on RNN, transformers, and conformers. It has an in-built data loader for MuST-C, Librispeech, and CoVoST datasets.
  • 3.
    NeurST [^11] https://github.com/bytedance/neurst (Zhao et al., 2020) is a lightweight toolkit, as it has no dependency on kaldi[CITE]. It has high computation efficiency using mixed precision and accelerated linear algebra and achieves faster training on large-scale datasets using Horovod and Byteps [CITE].
  • 4.
    SLT.KIT [^12] https://github.com/isl-mt/SLT.KIT(Zenkel et al., 2018) offers ASR, MT and ST models along with some specific features such as CTC and Attention based ASR, ASR with punctuation and a neural MT system.
Figure 1: End-to-end ST models taxonomy
5 End-to-End ST Models
End-to-end models for ST as discussed previously are gaining traction comparably to cascade models. This section presents an overview of E2E models as shown in Fig. 1. We categorize them under two major E2E themes. The first is based on the framework used, such as ASR/MT as pre-training objectives or modality-bridging technique. The second category is based on the nature of the data. We note that all of these methods rely on encoder-decoder (Seq2Seq) architecture for sequence learning (see Fig.2 (a)), where the encoder receives input from audio and/or text, while the decoder produces the translated text. The categorization presented here is based on which component boosts the ST task performance, as claimed in the papers. As such, the demarcation is not strict, and there may be overlaps in the categories. In addition, our emphasis in the present review of existing works is highlighting the core contribution and limitations as claimed by the authors. That means we look for answers to the question: what is the main technical contribution of authors to solve the ST problem? Thus, wherever possible, we have limited the mathematical description and believe that such details can be found in the related papers. Our attempt is to provide a succinct and clear picture of what works and what does not while addressing the ST problem.
(a) Framework based on Seq2Seq
(a) Framework based on Seq2Seq
(b) Framework based on Modality Bridging
Figure 2: E2E Framework-based ST Models. (a) Seq2Seq model, (b) Seq2Seq with modality bridging component. The dashed arrow denotes optional components.
5.1 ST Models based on Frameworks
As mentioned in the previous section, ST models in the literature employ the seq2seq framework alone or in combination with some modality-bridging component.
5.1.1 Seq2Seq Frameworks
A Seq2Seq model generates a sequence of outputs from a sequence of inputs as shown in Fig. 2(a). It has an encoder for speech input, a decoder for text output, and an optional shared/semantic decoder connecting the encoder and the decoder. The model is usually optimized for the ST loss. The Seq2Seq models have mainly two components that form the core of the models proposed: attention and transformer. This section discusses only the works which do not use external data for training. Models that leverage external data for training such as ASR and/or MT are discussed in the next section.
Attention mechanism is used to concentrate on specific sections of the input data for producing the output (Vaswani et al., 2017). It has been a successful strategy for getting state-of-the-art (SOTA) results in NLP as well as other areas. Below we provide efforts made to handle ST tasks using the attention mechanism within the Seq2Seq framework.
The author in (Bérard et al., 2016) uses a Seq2Seq model that eliminates the need for source transcript in the ST task. They use convolutional attention to “remember” and avoid translating the signal twice, which outperforms a hierarchical encoder with better results on synthetic data. The same author in (Bérard et al., 2018) uses source transcript and achieves results close to cascade models on LibriSpeech data. In (Duong et al., 2016), the author proposes phone-to-text alignment with a structural bias feature in the attention model. The measurement of alignment has been explored in (Anastasopoulos et al., 2016), which used IBM’s translation model as well as dynamic time warping131313dynamic time warping (DTW) is an algorithm for measuring similarity between two temporal sequences, which may vary in speed.. Seq2seq with attention trained using multitask learning achieves promising results in (Weiss et al., 2017). These models struggle with noisy inputs and long acoustic signals (Kim et al., 2017). They use a joint CTC-attention model trained through multitask learning by incorporating regularizers. The author uses two decoders where the second decoder seeks higher level representation (HLR) from the first decoder besides the encoder via the attention mechanism using a small dataset.
Transformer is the architecture based on multi-headed self-attention (Vaswani et al., 2017) which produces contextulized representation of the input. Because of parallelization and contextual representation, transformers have outperformed RNNs on several NLP tasks. This entails us applying transformers for the ST task as well.
Transformer-based Seq2Seq with attention is proposed in (Di Gangi et al., 2019). The architecture has a quadratic memory complexity, which involves: (a) CNN to downsample the input, and (b) 2-D attention to address short-range dependencies of spectrograms. In (Alastruey et al., 2022), the weight of some attention is avoided for speech tasks, hence decreasing the size of the attention matrix. The transformer encodes the speech features, thereby introducing local self-attention with a suitable window size for each layer to reduce the computational complexity.
In table 2, we present performance scores (BLUE/CER) of various ST models based on the Seq2Seq framework. We note that many of these results are not directly comparable since they use different amounts of speech data for testing.
Models/Techniques
Problem Solved
Dataset
Language Pair
Speech (hours)
Metric (BLEU)
Attention (Bérard et al., 2016)
Eliminate Source Transcription
BTEC (Synthetic)
[…]
15
46.7
Phone to word Alignment (Duong et al., 2016)
Word-spotting + low resources
CALLHOME
[…]
38
21.2
FastAlign + DTW (Anastasopoulos et al., 2016)
low resource
CALLHOME
[…]
38
30.8
Fisher
[…]
20minutes
53.8
Multitask Learning (Weiss et al., 2017)
performance
CALLHOME
[…]
38
48.7
Fisher
[…]
163
17.4
Joint CTC-Attention Model (Kim et al., 2017)
Noisy Signals, long inputs
WSJ1
[…]
80
(CER-7.36)
WSJ0
[…]
15
(CER-14.53)
Chime-4
[…]
18
(CER-44.99)
Encoder + Decoder (Bérard et al., 2018)
Cascading
Librispeech
[…]
236
22.2
BTEC (Synthetic)
[…]
15
50.7
Quadratic Memory Complexity (Di Gangi et al., 2019)
High training time
IWSLT
[…]
80
8.94
MuST-C
[…]
408
12.93
Drop weights that attention discards (Alastruey et al., 2022)
Attention for long sequence
MuST-C
[…]
408
22.49
[…]
504
27.46
[…]
465
22.97
Table 2: Performance scores of the models in Seq2Seq framework with the datasets, language pairs, duration of speech, and metric(BLEU).
Findings: Our study of Seq2Seq-based frameworks for ST task reveals that (a) structural bias can be obtained by stacked/pyramidal RNN and alignment smoothing, (b) regularizers such as transitivity and invertibility improves Character Error Rate, (c) HLR helps in transcription as well as translation, and (d) changing the self-attention of the encoder with a logarithmic distance penalty enhances translation.
ST with ASR: In this section, we delve into the research that leverages ASR data to pre-train a Seq2Seq model, subsequently fine-tuning the model for the ST translation task.
Curriculum pre-training (Wang et al., 2020d) refers to the use of ASR data for pre-training an Seq2Seq model, which is transformer-based, allowing it to learn transcription. The author argues that if the model is further pre-trained on learning semantic concepts (via frame-based masked language modeling) and word alignment (via frame-based bilingual lexical translation), it boosts the ST task performance. Specifically, existing E2E models either pre-train the encoder or use multi-task learning for ST tasks. As such, the encoder cannot isolate the learning of three tasks: transcription, semantic concept, and alignment, which are segregated by dividing the labor, and experiments prove the theoretical claims. Listen, Understand, and Translate (LUT) from (Dong et al., 2021a) uses the seq2seq model with external ASR data. Their primary contribution is to introduce a semantic encoder network, whose task is to use the encoder’s output from transcription to minimize the mean-squared loss between the semantic representations and the BERT embeddings of the target text. Such a strategy implicitly builds and trains an NMT model for translation. (Zhang et al., 2022) uses neural acoustic feature modeling (NAFM) to obtain the speech features from the audio instead of using Fbank, which leads to information loss.
The BLUE scores are given in Table 3 to provide readers with the quantitative scores of ST models employing ASR data.
Models/Techniques
Problem Solved
Dataset
Language Pair
Speech (hours)
Metric (BLEU)
Curriculum Pre-training + Decoder Pre-training (Wang et al., 2020d)
Heavy burden on Encoder
Librispeech
[…]
236
18.01
IWSLT
[…]
271
18.15
Listen Understand Translate (Dong et al., 2021a)
lack of parallel corpus
Librispeech
[…]
100
17.75
IWSLT
[…]
272
16.35
TED
[…]
524
20.84
NAFM + Parameterized Distance Penalty (Zhang et al., 2022)
lack of transcripts
MuST-C
[…]
452
25.2
CoVoST-2
[…]
427
11.7
[…]
427
17.3
Librispeech
[…]
100
18.9
Kosp2e
[…]
190
5.8
Table 3: Performance scores of the ST models with ASR showing the datasets, language pairs, duration of speech, and metric(BLEU).
ST using MT: This section discusses approaches that use either MT data for pre-training or directly using pre-trained MT in the ST decoder. These techniques rely on the idea of generating pseudotext and then translating them using MT. In (Bansal et al., 2017), the author uses the Unsupervised Term Discovery (UTD) approach to group repeated words into pseudo-text, which is subsequently used for training an MT model for bag-of-words using the parallel pseudo-text and target translations. The main advantage of such a system is that it can translate some content words under low-resource settings. The overall results are not very promising on the Spanish-English Call-Home dataset. Another limitation of this work is that the approach is not an E2E in a true sense as it involves two models– a UTD and then an MT model. A weakly supervised learning method for ST (Jia et al., 2019) that outperforms multi-task learning takes advantage of the MT and TTS synthesis models, which are pre-trained. In their study, (Liu et al., 2019) a transformer-based “teacher” model for the MT task and a “student” model for the ST task are trained via knowledge distillation (KD). They, however, rely on source language text and do not improve upon the pipeline system. Following along, (Gaido et al., 2020) explores word, sentence and sequence-interpolation based KD approaches for transferring knowledge from pre-trained MT to ST model. Empirical results of ST models trained using MT auxiliary data are presented in Table 4.
Models/Techniques
Problem Solved
Dataset
Language Pair
Speech (hours)
Metric (BLEU)
Unsupervised Term Discovery (Bansal et al., 2017)
low resource (no ASR & no MT)
CALLHOME
[…]
11
(Recall- 54.4)
Pre-trained ASR & MT (Jia et al., 2019)
lack of parallel corpus
Librispeech
[…]
26.7
Knowledge Distillation (Liu et al., 2019)
one model for SR & TT
Librispeech
[…]
100
22.91
MuST-C
[…]
542
19.55
KD with word-level/ seq.-level/ seq. interpolation (Gaido et al., 2020)
seq2seq for ASR & MT
Librispeech
[…]
100
16.8
Table 4: Performance scores of the ST models with MT showing the datasets, language pairs, duration of speech, and metric(BLEU).
Findings: (a) UTD suffers from low recall and precision, and (b) transferring knowledge via KD from pre-trained MT to ST causes gender bias, omission of sentences, and generic verbal-tense choice.
ST using both MT and ASR: This section elucidates the work that employs both MT and ASR data for ST tasks.
In (Bahar et al., 2019), the author presents a comparison of various E2E ST architectures such as direct, multitask many-to-one, one-to-many, tied-cascade, and tied-triangle and shows faster convergence when the CTC loss is combined with other losses. They show that pre-trained models with ASR and MT losses achieve promising results. Contrary to claims of (Anastasopoulos and Chiang, 2018), tied-triangle architecture is no better than a direct model when fine-tuned properly. Since the ST task is similar to the MT task from the output perspective, works such as XSTNet (Ye et al., 2021) utilize external MT data to pre-train the encoder-decoder network extensively, then fine-tune it using parallel corpus data of MT, ST, ASR, and external MT data for optimizing the model using what they call progressive training. They achieve impressive performance on MuST-C and augmented Librispeech data. They also demonstrate improved performance on auxiliary tasks of MT and ASR. STPT model (Tang et al., 2022) proposes four sub-tasks for multitask pre-training: text-to-text (T2T), which is self-supervised; speech-to-phoneme which is supervised; acoustic learning, which is self-supervised, and ST which is supervised. Only T2T and ST tasks would subsequently be used for fine-tuning. Despite pre-training on “unlabeled” speech data, they obtained superior results on MuST-C data for the ST task. Table 5 shows BLUE scores of models utilizing both ASR and MT data.
Models/Techniques
Problem Solved
Dataset
Language Pair
Speech (hours)
Metric (BLEU)
Tied multi-task learning with regularizers (Anastasopoulos and Chiang, 2018)
low resource on speech transcription & translation
CALLHOME
[…]
20
28.8
Glossed Audio Corpus [^14] https://ainu.ninjal.ac.jp/folklore/en/
[…]
2.5
20.3
Godard Corpus [^15] https://github.com/besacier/mboshi-french-parallel-corpus
[…]
4.4
24.7
CTC loss (Bahar et al., 2019)
unknown segmentation of input sequence
IWSLT
[…]
390
24.43
Librispeech
[…]
200
15.74
Cross Speech-Net Network (XSTNet) (Ye et al., 2021)
Error Propogation
MuST-C
[…]
408
27.8
[…]
504
30.8
[…]
492
38.0
[…]
465
26.4
[…]
442
31.2
[…]
385
32.4
[…]
432
25.7
[…]
489
18.5
Librispeech
[…]
200
21.5
Speech & Text Joint Pre-training (Tang et al., 2022)
subtask interference
MuST-C
[…]
504
33.1
[…]
492
39.7
Table 5: Performance of the ST models with both ASR & MT showing the datasets, language pairs, duration of speech, and metric(BLEU).
Findings: (a) Progressive training needs a huge data and training time in order to achieve superior results, (b) multitask pre-training can be used to leverage unlabeled speech data.
Other: In (Sperber et al., 2019), the author analyzes that the cascade ST models, which use ASR and MT models, require less data than the direct ST models. But the drawback of the cascade two-stage ST models is to propagate errors from ASR to MT models and thus to ST model. Therefore, the author introduces an Attention-Passing Model (APM), which only passes high-attention vectors from the audio encoder to the translation text for decoding.
Findings: APM requires less data than cascade ST models.
5.1.2 Seq2Seq with Modality Bridging
A method for learning a combined representation of text and speech is called modality bridging (see fig.2 (b)). Both speech and text in an ST task represent the same semantic meaning. Hence, a good ST model should learn a representation such that embeddings of both modalities for similar speech-text pairs lie close to each other. It is believed that low performance on ST tasks is due to models not learning aligned representations of speech and text. Therefore, different authors have devised different ways to fill the gap, which fall into five major approaches: (a) adapters, (b) contrastive learning, (c) Knowledge-distillation, (d) optimal transport, and (e) mix-up strategy. Below we discuss the works utilizing these approaches and show the pros and cons.
Adapters are small modules integrated with pre-trained networks for specific tasks (Houlsby et al., 2019). They have performed at par with fine-turning-based approaches while requiring only a fraction of trainable parameters. For example, in (Zhao et al., 2022; Gállego et al., 2021), the modality gap is filled using adapter layers, which is a multi-headed self-attention with pooling operation. The author uses Wave2Vec 2.0 (Baevski et al., 2020) for speech-feature extraction, wherein self-attention layers in the transformer are equipped with pooling operation for dimensionality reduction to match text representation.
Contrastive learning approximates the “semantic” distance in the input space using a simple distance in the target space after mapping input patterns onto the target space (Chopra et al., 2005). It tries to bring positive instances closer while pushing negative ones apart. It has been used excessively in both supervised and unsupervised settings for learning representations. Below, we provide details of works leveraging contrastive learning loss for modality bridging.
The author in (Zhang et al., 2023) performs the explicit knowledge transfer through contrastive learning. They learn frame and sentence-level speech feature representation and use whitening (Su et al., 2021) to alleviate the MT representation degeneration. In (Liu et al., 2019), they decouple the encoder representation into three parts: acoustic encoder, shrinking (done via CTC) of acoustic encoder output, and semantic encoder for modality-gap bridging. Using a contrastive learning architecture, Chimera (Han et al., 2021) trains a semantic memory module which is shared for overcoming the modality distance. XSTNet (Ye et al., 2021) augmented with contrastive loss (Ye et al., 2022a) investigates three different methods: span masked representation, word-repetition and cut-off. It is also found that contrastive loss is better than CTC and L2 loss. The same author in (Ouyang et al., 2022) proposes word-aligned contrastive learning (WACO) by forming average speech and word embedding of the same word as positive pair while of different words as negative pairs. CSTNet is a self-supervised learning framework based on contrastive learning (using a mix of triplet losses)(Khurana et al., 2020). On top of the CTC loss, the boundary-based speech length shrinking mechanism is applied in (Zeng et al., 2022). The authors claim that if boundary-based shrinking is applied with other modality-bridging techniques, such as contrastive loss, it can further improve the model performance. The approach presented achieves lower inference speed and memory footprint.
Knowledge-distillation (Hinton et al., 2015) is a mechanism to distill information from a trained and large “teacher” model to a smaller and efficient “student” model. It has been used with $L_{2}$ loss in (Huzaifah and Kukanov, 2023) to address the modality gap issue.
Optimal transport (Peyré et al., 2019) is a mechanism for comparing two probability distributions . In the ST task, speech and text representations may be deemed as two probability distributions. Recently, (Le et al., 2023) used optimal transport (OT) and CTC together in order to close the modality gap during the pre-training phase. They show significant gains in BLUE score when the ST model is fine-tuned without any external data compared to multitask learning.
Mix-up strategy Speech-Text Manifold Mixup (STEMM) (Fang et al., 2022) strategy uses speech embedding and mixes embeddings of speech and text into encoder-decoder of a translation model for bridging the modality gap under the self-supervised learning framework.
Models/Techniques
Problem Solved
Dataset
Language Pair
Speech (hours)
Metric (BLEU)
Knowledge Distillation (Liu et al., 2019)
performance of text translation
Librispeech
[…]
100
17.02
MuST-C
[…]
542
19.55
M-Adapter + W2V2 + mBart (Baevski et al., 2020)
training gap between Pre-training & Fine-tuning the modality
MuST-C
[…]
408
25.9
[…]
432
24.62
[…]
492
37.34
Chimera (Han et al., 2021)
projecting audio & text to a common semantic representation
MuST-C
[…]
408
27.1
[…]
492
35.6
[…]
489
17.4
[…]
504
30.6
[…]
465
25.0
[…]
432
24.0
[…]
385
30.2
[…]
442
29.2
Librispeech
[…]
492
19.4
ConST (XSTNet + Constrastive Loss) (Ye et al., 2021)
closes modality gap
MuST-C
[…]
408
28.3
[…]
504
32.0
[…]
492
38.3
[…]
465
27.2
[…]
442
31.7
[…]
385
33.1
[…]
432
25.6
[…]
489
18.9
W2V2 + mBart + Adapter (Gállego et al., 2021; Zhao et al., 2022)
slow convergence speed
MuST-C
[…]
408
28.22
IWSLT
[…]
390
21.43
WACO (Ouyang et al., 2022)
limited parallel data (1-hour)
MuST-C
[…]
1
17.5
IWSLT
[…]
1
13.3
AdaTrans (Zeng et al., 2022)
closing gap between length of speech & text
MuST-C
[…]
408
28.7
[…]
492
38.7
[…]
489
19.0
STEMM (Fang et al., 2022)
Speech representation
MuST-C
[…]
408
28.7
[…]
492
37.4
[…]
489
17.8
[…]
504
31.0
[…]
465
25.8
[…]
432
24.5
[…]
385
31.7
[…]
442
30.5
CTC loss + Optimal Transport (Siamese-PT) (Le et al., 2023)
without change in architecture
MuST-C
[…]
408
27.9
[…]
504
31.8
[…]
492
39.2
[…]
465
27.7
[…]
442
31.7
[…]
385
34.2
[…]
432
27.0
[…]
489
18.5
CoVoST-2
[…]
21.5
[…]
25.5
Fine & Coarse Granularity Contrastive Learning (Zhang et al., 2023)
limited knowledge transfer ability
MuST-C
[…]
408
25.9
[…]
492
38.3
[…]
489
19.7
[…]
504
31.9
[…]
465
27.3
[…]
432
26.8
[…]
385
32.7
[…]
442
31.6
Table 6: Performance of the ST models using modality bridging. The datasets, language pairs, duration of speech, and metric(BLEU) are shown.
Table 6 presents the performance scores of ST models based on modality-bridging techniques. We can observe that mixup strategy achieves the highest BLUE score on En-De pair. Whereas boundary-based speech length shrinking mechanism matches the score when combined with other modality-bridging techniques.
Findings: (a) adapters can shrink the speech length as well as the modality distance between the text and speech representations while requiring a small number of trainable parameters, (b) Contrastive loss is better than CTC and $L_{2}$ loss for modality-bridging, (c) boundary-based speech length shrinking combined with contrastive loss can improve the ST task performance.
5.2 ST Models based on the Nature of Available Data
In the previous section, we provided an overview of the ST models based on the frameworks used. The present section provides readers with another perspective on E2E ST models. In particular, it discusses the E2E ST models categorized based on the nature of the data, such as data is low-resource, streaming, multilingual, etc. Given the specific challenges they pose, we believe such a categorization might be interesting to researchers.
5.2.1 ST in Low-Resource settings
A low-resource language (LRL) is one where speech and/or text data is available in smaller quantity– usually not enough to pre-train Seq2Seq models. As such, LRLs present challenges of their own such as overfitting and poor generalization. This section will discuss works where ST models are developed especially for low-resource languages. The proposed models under this category have generic architecture as shown in Fig.3a which is similar to Seq2Seq ST models. We find the approaches mainly use pre-training the encoder on high-resource ASR data and subsequent fine-tuning on ST data. For example, (Bansal et al., 2018) empirically demonstrates 100% performance improvement on ST tasks. They find that if the ASR language differs from the source and target languages, then pre-training on ASR data enhances ST task performance. Though the BLUE score is improved, the absolute BLUE score is only 7.1. In (Wang et al., 2022a), the unsupervised ST is implemented for low-resource settings using pseudo-labels from unsupervised cascade models. Table 7 presents scores of LRL ST models.
Findings: Low-resource ST can be tackled via either pre-train on ASR data, and then fine-tuning on ST data or using pseudo-labels.
Models/Techniques
Problem Solved
Dataset
Language Pair
Speech (hours)
Metric (BLEU)
Encoder-Decoder + Attention (Bansal et al., 2018)
Low- resource
Fisher
[…]
20
20.2
Godard Corpus
[…]
4
7.1
Unsupervised subtasks + W2V2 + mBart (Wang et al., 2022a)
Pseudo-labels & low downstream performance
CoVoST
[…]
264
40.1
[…]
113
43.8
[…]
16
48.6
[…]
3
19.0
[…]
2
25.0
[…]
38.5
[…]
22.2
[…]
22.1
Table 7: performance of ST models on low resource languages with the datasets, language pairs, duration of speech, and metric(BLEU).
Figure 3: E2E ST Models based on the nature of the data. (a) Low-resource ST, (b) Code-Mix ST, (c) Wait-K based Streaming ST, (d) RNN-T based Streaming ST, (e) Unsupervised ST, and (f) Multilingual ST. The dashed arrow denotes optional components.
Figure 4: Wait-k strategy for streaming ST setting. In this, the decoder waits for $k$ input speech segments before starting to output. Thereafter, it produces one token for every source segment. The figure showcases the scenario with $k=2$.
5.2.2 Code-Mix ST
Code-Mix language refers to speech where one primary language is used, but words or phrases from other (embedded) languages are also included. We find that there exist only a few works on code-mix ST. In (Weller et al., 2022), the code-mix dataset is created with the existing publicly available corpora Fisher (Cieri et al., 2004) and Miami [^16] https://github.com/apple/ml-code-switched-speech-translation. As shown in Fig. 3b, Code-mix ST models feed language ID in addition to speech input to the encoder of the Seq2Seq model (Weller et al., 2022). The Wav2Vec 2.0, an acoustic encoder, and mBART, a multilingual decoder, are used for both languages with an attention layer applied for the embedded language. The results are shown in Table 8.
Findings: Standard ST models achieve good results on code-mix data with low-resource setting and no fine-tuning, especially when encoders and decoders are pre-trained, and multilingual models are used.
Models/Techniques
Problem Solved
Dataset
Language Pair
Speech (hours)
Metric (BLEU)
Joint Transcription & Translation + LID (Weller et al., 2022)
Mixed Language Speech
Fisher
[…]
15
26.2
Miami
[…]
13
18.3
Table 8: Performance of ST code-mix Models: Datasets, Language Pairs, Speech Duration, and BLEU Metric.
5.2.3 ST in Streaming Setting
ST in the streaming setting is the task of simultaneous translation (SST) (Goldman-Eisler, 1972; Fügen et al., 2007; Grissom II et al., 2014), that is, the translation of the input as soon as it arrives without waiting for the entire input. It finds application in online speech translation and video dubbing, inter alia. Traditionally, the streaming ST problem has been solved by feeding the segmented output of a streaming ASR model to a streaming MT model (Iranzo-Sánchez et al., 2020; Oda et al., 2014). However, due to the cascade nature of the model, it is prone to high latency and error propagation.
The generic architecture of the streaming ST model is shown in Fig. 3c and 3d. The streaming or simultaneous ST problem has been approached in broadly the following ways: Wait-K, Segmentation, Integrate-and-Fire (IF), and transducers/unidirectional transformers. These techniques are combined with Seq2Seq frameworks to devise streaming ST models.
Wait-K policy (Ma et al., 2018) learns the parameters $\theta$ of the model by optimizing the negative log-likelihood $-\sum_{(\mathbf{u,v})\in D}\log p(\mathbf{v}|\mathbf{u};k;\theta)$, where $k$ is the number of segments to look before starting translation (see Fig. 4). The probability $p(\cdot)$ is calculated as
$p({\bf v}|{\bf u};k;\theta)=\prod_{t=1}^{T_{v}}p(v_{t}|v_{<t},u_{t+k};\theta)$ (21)
Wait-K policy guarantees that the model can look at $t+k-1$ speech segments while predicting token $y_{t}$ (Ren et al., 2020). However, one limitation of the Wait-K policy is that it fails to do a beam search while decoding except for long-tail (Ma et al., 2018). To solve this problem, (Zeng et al., 2021) proposes a Wait-K stride-N policy. It essentially is a Wait-K policy with the addition of N READ and WRITE operations until the end of the sentence after reading the first K-segments. In order to determine the K-segments, (Chen et al., 2021b) leverages streaming ASR to guide the direct simultaneous ST decoding via beam search.
As discussed above, determining when to write is crucial for efficient SST. Contrary to Wait-K policy, Segmentation can be performed on the embedded speech using either an incremental BEAM search (Yan et al., 2023) or using attention mechanism (Papi et al., 2022). Essentially, these works adapt offline ST to SST showing spectacular performance on benchmark datasets. Note that the models proposed in (Yan et al., 2023; Papi et al., 2022) train models in a cascade manner while the inference is E2E.
The aforementioned works require that encoded speech be segmented so that the decoder can apply the Wait-K policy. The goal of segmentation is to identify the word, sub-word, or phone boundary which are usually not even (due to silences, longer syllables, etc.). That means the number of acoustic units varies with time in each segments. Monotonic-segmented Streaming ST (MoSST) (Dong et al., 2021b) is based on learning when to translate, which has a monotonic segmentation module located between the acoustic encoder and the transformer. It has an Integrate-and-Fire (IF) neuron (Abbott, 1999), which fires above a threshold when the context is developed. If the context is not developed, the neuron receives signals and keeps accumulating the acoustic vectors. IF strategy has shown impressive performance in simultaneous ASR (Dong and Xu, 2019) and ST (Chang and yi Lee, 2022). It can be used for monotonic segmentation of the speech input along with adaptive decision strategy (Dong et al., 2021b).
RNN-T are transducers that can output text sequences given the input speech sequence in an online/streaming fashion (Graves, 2012). For example, Cross-Attention Augmented Transducer (CAAT) optimizes translation and policy models in tandem (Liu et al., 2021). They eliminate the RNN-T’s strict monotonic restriction for reordering in the translation. The use of transformers to reduce the multi-step memory footprint causes a significant delay for CAAT. The use of regularization terms and substantial hyperparameter adjustment are some limitations of CAAT. An extension of it in (Xue et al., 2022) leverages Transformer Transducer (TT) networks with attention pooling for streaming E2E ST tasks. Attention divides the input audio into chunks of specific sizes. At any time, processing any input frame $\bf x_{t}$ can only see frames within its own chunk and a fixed number of left chunks. By sharing the encoder, they also propose a variant to handle E2E ST tasks in multilingual settings. The adaptive READ and WRITE policy choices between encoder output and ground truth contributed to its success. The same authors (Wang et al., 2022b) propose to combine the benefits of language-specific and language-agnostic encoders within the TT framework. A shared encoder takes LIDs as gating values and computes weights for each language through the source LID scheduling scheme. The empirical results demonstrate superior performance and lesser trainable parameters compared to bilingual ST. The results of SST models are shown in Table 9.
Models/Techniques
Problem Solved
Dataset
Language Pair
Speech (hours)
Metric (BLEU)
SimulSpeech + Attention-level KD + Data-level KD (Ren et al., 2020)
online streaming setting
MuST-C
[…]
496
22.49
MoSST (Dong et al., 2021b)
finding boundaries for acoustic units
MuST-C
[…]
408
24.9
[…]
492
35.3
CAAT (Liu et al., 2021)
policy & model for reordering in translation
MuST-C
[…]
408
22.3
Transformer Transducer(TT) (Xue et al., 2022)
High Inference latency & error propagation
MSLT Corpus
[…]
50K
30.7
[…]
50K
35.6
LAMASSU (Wang et al., 2022b)
Joint ASR & ST Model
(closed dataset)
[…]
10K
23.6
[…]
10K
25.5
[…]
10K
23.3
Table 9: Performance of the ST models for Streaming (Simultaneous Translation) data: Datasets, language pairs, duration of speech, and Metric(BLEU).
Findings: (a) RNN-T are able to reduce memory footprint at the expense of translation delay, (b) adaptive read-write policy improves ST task performance.
5.2.4 Unsupervised ST
There is an abundance of unlabeled speech and text data. Since manual annotation and creating a parallel corpus is costly, the natural instinct is to exploit unlabeled data for training ST models. This section reviews works where researchers make use of the unlabeled speech data to advance the ST task performance.
In (Wang et al., 2021b), the author leverages large-scale self- and semi-supervised learning. They pre-train a Wave2Vec 2.0 as an encoder on Libri-light data (Kahn et al., 2019), and the decoder is randomly initialized. The entire model is optimized on CoVoST 2 ST data, and the encoder is frozen. Thereby, self-training is executed to generate pseudo-labels for Libri-light data. The Wav2Vec 2.0 is a “student” model which is fine-tuned with ground truth CoVoST 2 data and pseudo labels. Finally, a language model (LM) is trained on CommonCrawl data and combined with the ST model to generate text via beam-search decoding. Following along, for training E2E model, (Wang et al., 2021b) produces pseudo-labels by cascading ASR, text de-normalization, and MT in an Unsupervised manner. In (Li et al., 2020), Wave2Vec 2.0 and mBART are optimized for domain adaption using in-domain data. According to experimental results, the proposed method is effective for E2E models without pre-training. However, between supervised and unsupervised pre-trained models performance gap is encountered, which may be investigated in future works. The results of applying unsupervised ST models are presented in Table 10.
Models/Techniques
Problem Solved
Dataset
Language Pair
Speech (hours)
Metric (BLEU)
Teacher-Student Model (W2V2 + self-training + dec. w/o LM) (Kahn et al., 2019)
Unlabelled Speech & Data
CoVoST-2
[…]
430
27.2
[…]
430
35.6
[…]
430
20.8
[…]
430
18.9
Table 10: Performance of ST models trained in an unsupervised setting: Datasets, Language Pairs, Speech Duration, and BLEU Score Analysis.
Findings: (a) Pre-trained acoustic and language models combined with pseudo-labels using self-training perform well for unsupervised ST translation, (b) domain adaptation is another approach for unsupervised ST.
5.2.5 Multilingual ST
The multilingual ST model aims to translate from/to multiple speech input/output languages. It can be one of many-to-one, one-to-many, or many-to-many. The ST models solve multilinguality issues using mainly three approaches: (a) language ID, (b) dual-decoder, and (c) pre-trained models.
Language ID (LID) is the identification label that allows one to identify the target language and explicitly translate the speech simultaneously. The existing works handle multilinguality using LID either with encoder or decoder. In (Inaguma et al., 2019), the model uses LID in the decoder for one-to-many and many-to-many translation. They demonstrate impressive performance in translation from high-resource to low-resource languages without using any transcript data from LRL. However, using the LID embedding in the decoder (Gangi et al., 2019) is shown to underperform than using it in the encoder. The author shows that LID can be either concatenated or merged with the inputs and, when pre-trained with ASR data, can result in superior performance than the one-to-one system. The model, however, performs poorly when trained on many unrelated target languages. One-to-many and many-to-one multilingual ST systems of (Wang et al., 2020c, a) provide a good set of baselines for research purposes.
Dual-decoder model is the transformer with two decoders, one for each ASR and ST, and the dual-attention mechanism. In (Le et al., 2020), a dual-decoder model is proposed to optimize it for ASR and ST tasks jointly. The author hypothesizes that a dual-attention mechanism can benefit each task by transferring knowledge instantly or in wait-k policy mechanism. Their model generalizes earlier models proposed for one-to-many and bilingual ST models.
Pre-trained Multilingual Models use a pre-trained encoder and decoder for acoustic modeling and language modeling, respectively. In (Li et al., 2020; Tran et al., 2020), the author shows that efficiently fine-tuning mBART, which is a pre-trained multilingual decoder (Liu et al., 2020a) can achieve SOTA results on CoVoST data on zero-shot cross-lingual and multilingual translation tasks. Along similar lines, (Le et al., 2021) shows that inserting adapters in between layers of the encoder-decoder framework and tuning them can improve the ST task performance over bilingual ST models. The comparative performance statistics of multilingual ST models are presented in Table 11.
Models/Techniques
Problem Solved
Dataset
Language Pair
Speech (hours)
Metric (BLEU)
LID + Mixed Data Training (Inaguma et al., 2019)
one-to-many & Many-to-many ST
Fisher
[…]
163
46.3
CALLHOME
[…]
20
17.3
Librispeech
[…]
236
17.6
ST-TED
[…]
18.6
Dual-Decoder(Esp-Net) (Le et al., 2020)
Joint ASR & ST
MuST-C
[…]
408
23.63
[…]
504
28.12
[…]
492
33.45
[…]
465
24.18
[…]
442
27.55
[…]
385
29.95
[…]
432
22.87
[…]
489
15.21
LNA finetuning + Zero Shot (Liu et al., 2020a)
unseen languages & low resource
CoVoST-2
[…]
264
36.1
[…]
184
30.8
[…]
113
38.1
[…]
136
31.8
[…]
44
31.9
[…]
18
30.9
[…]
10
24.1
[…]
7
25.7
[…]
2
5.6
[…]
2
5.9
[…]
49
11.0
[…]
10
9.1
[…]
4
11.2
[…]
3
2.9
[…]
3
1.2
[…]
2
6.4
[…]
2
5.0
[…]
2
9.0
[…]
2
0.9
[…]
1
2.1
[…]
1
3.7
[…]
18.6
[…]
30.9
[…]
30.6
[…]
26.6
[…]
22.2
[…]
21.5
[…]
29.9
[…]
39.3
[…]
21.5
[…]
14.8
[…]
25.1
[…]
30.4
[…]
17.8
[…]
17.0
[…]
34.4
Adapters (Le et al., 2021)
Performance
MuST-C
[…]
408
24.63
[…]
504
28.73
[…]
492
34.98
[…]
465
24.96
[…]
442
28.8
[…]
385
31.13
[…]
432
23.78
[…]
489
16.36
Table 11: Comparison of multilingual ST models performance Using Datasets, Language Pairs, Speech Length, and the BLEU Metric.
Findings: (a) Language ID token works well with the encoder, (b) Mixed data training with language ID can transfer learning from HRL to LRL, (c) LID training with ASR data on unrelated languages does not work well, (d) adapters can help boost multilingual pre-trained models translation performance.
5.3 Discussion
The works presented so far show that E2E ST models have been improved tremendously. ST models’ improved performance is likely due to leveraging pre-trained ASR/MT models or the respective corpus to train ST encoders/decoders. Weakly labeled/pseudo labels are another way to create more data for training ST models. Contrastive learning, mix-up strategy, adapters, and optimal transport are a few ways to bridge the modality gap.
Applying unsupervised ASR and MT with the Wav2Vec 2.0 encoder and mBART decoder in a low-resource setting yields good results for ST models. When considering online data streaming, using the IF neuron for context building and translation improves results compared to using CAAT, which had latency issues due to reordering for translation tasks introduced by RNN-T. mBART handles multilingual settings well by using a dual attention mechanism that facilitates knowledge transfer. Additionally, inserting adapters between the encoder and decoder layers improves performance. In the unsupervised ST setting, the SOTA results were achieved by training Wav2Vec 2.0 on data within the same domain as the speech. We see that k-wait policy is used in the streaming settings with segmentation and Multilingual settings with a dual-attention mechanism. In both cases, it yields good results. Also, adapters are used in modality bridging and multilingual settings with pre-trained models, which improves the performance. Standard ST models perform satisfactorily on Code-Mix data requiring no fine-tuning. However, there has been limited study on Code-Mix data.
5.4 Overall Performance Trend of E2E ST approaches in Common Benchmarks
In this section, we analyze the performance evolution of ST models based on the MuST-C dataset, as depicted in Figure 5. We selected the MuST-C dataset due to its widespread adoption by researchers since its introduction in 2019. Figure 5 reveals a consistent enhancement in the performance of ST models over time. A noteworthy breakthrough is observed in June 2021, attributed to the work of Ye et al. (2021). This particular model achieved superior results by incorporating various key factors, including the utilization of Wav2Vec 2.0 for feature representation in place of fBank, the integration of a pre-trained MT decoder trained on extensive parallel MT data, and the application of multi-task fine-tuning across ASR, MT, and ST data.
Figure 5: Performance of ST models on MuST-C data over a period of three years.
Language Pair
Model/Technique
Dataset
Speech (hours)
Setting
Metric (BLEU)
[…]
Transformer Transducer (Xue et al., 2022)
MSLT Corpus
50K
Online Streaming
30.7
[…]
35.6
[…]
Speech & Text Joint Pre-Training (Tang et al., 2022)
MuST-C
492
ST with ASR & MT
39.7
[…]
Unsupervised subtasks + W2V2 + mBart (Wang et al., 2022a)
CoVoST-2
Low-Resource
38.5
[…]
22.2
[…]
16
48.6
[…]
CTC loss + Optimal Transport (Siamese-Pt) (Le et al., 2023)
MuST-C
465
Modality bridging
27.7
[…]
385
34.2
[…]
432
27.0
[…]
ConST & Siamese-Pt (Ye et al., 2021)
MuST-C
442
Modality Bridging
31.7
[…]
Multitask learning (Kim et al., 2017)
CALLHOME
38
Sequence to Sequence
48.7
[…]
Attention (Bérard et al., 2016)
BTEC
15
Sequence to Sequence
46.7
[…]
NAFM (Zhang et al., 2022)
Kosp2e
190
ST with ASR
5.8
[…]
LAMASSU (Wang et al., 2022b)
(closed dataset)
10K
Online streaming
23.3
[…]
LNA + Zero-Shot Learning (Li et al., 2020)
CoVoST-2
136
Multilingual
31.8
[…]
10
24.1
[…]
7
25.7
[…]
10
9.1
Table 12: SOTA performance in High-Resource Language Pairs: Dataset, Models, Speech Duration, Settings, and BLEU Score Analysis (- means the data is not mentioned explicitly in the paper)
Language Pair
Model/Technique
Dataset
Speech (hours)
Setting
Metric (BLEU)
[…]
FastAlign + DTW (Anastasopoulos et al., 2016)
Fisher
20minutes
Sequence to Sequence framework
(Fscore:53.8)
[…]
Tied Multitask Learning with regularizers (Anastasopoulos and Chiang, 2018)
Glossed Audio Corpus
2.5
ST with ASR & MT
20.3
[…]
Godard Corpus
4.4
24.7
[…]
WACO (Ouyang et al., 2022)
IWSLT
1
Modality Bridging
13.3
[…]
Unsupervised + W2V2 + mBart (Wang et al., 2022a)
CoVoST-2
3
Low-Resource
19.0
[…]
2
25.0
[…]
Teacher-Student (W2V2 + self-training + dec w/o LM) (Kahn et al., 2019)
CoVoST-2
430
Unsupervised
20.8
[…]
35.6
[…]
18.9
[…]
LNA + Zero Shot Learning (Li et al., 2020)
CoVoST-2
2
Multi-Lingual
5.6
[…]
2
5.9
[…]
49
11.0
[…]
4
11.2
[…]
3
1.2
[…]
2
6.4
[…]
2
9.0
[…]
2
0.9
[…]
1
2.1
[…]
1
3.7
[…]
430
30.6
[…]
430
22.2
[…]
430
21.5
[…]
430
29.9
[…]
430
39.3
[…]
430
21.5
[…]
430
14.8
[…]
430
25.1
[…]
430
30.4
[…]
430
17.8
Table 13: SOTA performance in Low-Resource Language Pairs: Dataset, Models, Speech Duration, Settings, and BLEU Score Analysis
5.5 SOTA Performance on High and Low-Resource Languages
In Table 12 and 13, we have compiled the SOTA performance of the ST models on high and low-resource language pairs, respectively, available at the time of writing which is Sep 2023. The tables show which models following which techniques are able to obtain the SOTA performance. Such a table gives a bird-eye view as well as where the performance of ST models stands today. We can observe from Table 12 that there are a few HRLs such as Ko and Zh where the performance of ST models is poor. Similarly, the results in Table 13 showcase really small BLUE scores for languages such as Mn, Si, Ta, Id, Ja, and Sv. Therefore, there is an immediate need to improve the performance of ST models for these languages.
6 Future Directions for Research
This section highlights challenges that need the attention of researchers working on ST problems.
6.1 Cascade vs End-to-End Models
As argued and presented through comprehensive experiments by (Bentivogli et al., 2021), the performance gaps between cascade and E2E ST models are bridged. However, as shown by (Agrawal et al., 2023) in a recent IWSLT 2023 subtitling generation task, the performance of cascade models is far superior to E2E models for offline ST tasks evaluated on all metrics. Furthermore, as far as our understanding, no thorough assessment has been done for low-resource languages that use E2E and cascade models. It may be interesting to compare E2E and cascade ST models on various ST datasets.
6.2 ST on Code-Mix data
We find that there is only one study on the ST model that uses code-mix data as an input. A code-mix data has problems, such as different lexicons, syntax, and scarcity of labeled data. Therefore, it will be interesting to (a) create Code-Mix ST datasets incorporating more languages, (b) see how the existing ST models perform on code-mix ST data?, and (c) Can pre-training in many languages assist in tackling the code-mixing issue.
6.3 Domain-Invariant Models
ST models developed for one domain do not scale well to other domains, as shown in the recent IWSLT 2023. Here domain in-variance setting is the ST model which is trained in some language combination (say Eng-De) and needs to be adapted to other language combinations (e.g., Eng-Hi). Transfer learning/continual learning can be explored to develop generic models.
6.4 Discrepancy between Automatic and Human Evaluation
There may be discrepancies and disagreements among various metrics used to report ST task results. They do not match the mean option score (MOS) provided by human evaluators (Agrawal et al., 2023). For example, if a system evaluates the BLUE score between a ground truth sentence “Police shot the culprit with a gun” and hypothesis sentence “Police use a gun to shot the culprit”, it is 0! However, both sentences above might be deemed appropriate translations of an utterance semantically by an ST system. Such an argument is supported by dubbing artists who often change the voice of the sentence to simplify it or make it more pleasing.  [^17] In the movie “Pirates of the Caribbean”, Jack Sparrow asks Bloom how long he can go for the girl. The original answer from Bloom is “I can die for her!”. Whereas Hindi dubbing is “Till the dying breadth”
Therefore, we call for the attention of researchers to develop and use metrics that match human evaluations semantically. An approach could be to subject the ground truth and hypothesis sentences under semantic textual similarity tasks and score them accordingly.
6.5 Handling Ambient Noise
In our literature survey, we find that little has been done to deal with ambient noises. Ambient noise, background music, cross-talk, and non-verbal sounds may create difficulty in ST model learning. The model must distinguish between a meaningful utterance and ambient noise– a non-trivial task.
6.6 Handling Multiple Speakers
It is common in the real world where the audio/video has multiple speakers, each of which may have its own accent (cf., An Asian and American talking to each other in English), dialect, pitch, and accent. Performing speech separation may be useful before feeding it to the ST model for improved performance.
6.7 Handling Speaker Diarization
Speaker diarization refers to demarcating the timing of speakers in a multiple-speaker speech. So far, the datasets for ST do not have speaker boundary marks. Creating speaker-diarized ST data in a multilingual setting will be interesting to test the ST models’ robustness.
6.8 Multilingual and Simultaneous ST
Multilingual ST has gained momentum recently due to its importance in the real world. For example, a single speech must be broadcast to multilingual communities (e.g., a conference is attended by a diverse group of people). It can be one-to-many, many-to-one, and many-to-many languages ST. Our literature survey shows that only a few works exist to target one-to-many and many-to-one, whereas many-to-many is few (Inaguma et al., 2019). Besides, there is an opportunity to explore simultaneous multilingual ST, which is the most practical setting.
6.9 Low-resource ST Datasets and Models
Most existing works have focused on building ST models and datasets for high-resource languages. As we know, the success of ST models relies on the parallel speech-text corpus; building ST datasets for low-resource languages requires immediate attention. Further, a few works, such as (Bansal et al., 2018), have reported ST task results on the Mboshi-French pair; however, the BLUE score is poor. Therefore, building models that transfer learning from language pairs with high to low resources is warranted.
7 Conclusion
This survey paper delves into the most recent advancements in E2E ST translation works. Our discussion includes models, evaluation metrics, and datasets used to train ST models. We review various frameworks for ST models and highlight previous research in this field. The categorization of ST models is based on the kind of data they handle and the models employed. Additionally, we discuss potential future directions for improving speech-to-text translation. Our findings suggest that the gap between cascade and E2E system performance in both online and offline settings is narrowing. However, for some language pairs, the gap is still wide and therefore, additional work is warranted. Our goal in the present ST survey is to offer valuable insight into this topic and drive advancements in ST research. We believe that such reviews will be interesting to researchers.